In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?
How would the equation look and how can I solve it?

Answer by robertb(5830)

1,800L + 2,400S = 48,000
L + S = 21.
You can use the method of substitution, letting L = 21 - S.